### v1

# Engineering RAG Systems for Real-World Applications: Design, Development, and Evaluation

Abstract
Retrieval-Augmented Generation (RAG) systems are emerging as a key approach for grounding Large Language Models (LLMs) in external knowledge, addressing limitations in factual accuracy and contextual relevance. However, there is a lack of empirical studies that report on the development of RAG-based implementations grounded in real-world use cases, evaluated through general user involvement, and accompanied by systematic documentation of lessons learned. This paper presents five domain-specific RAG applications developed for real-world scenarios across governance, cybersecurity, agriculture, industrial research, and medical diagnostics. Each system incorporates multilingual OCR, semantic retrieval via vector embeddings, and domain-adapted LLMs, deployed through local servers or cloud APIs to meet distinct user needs. A web-based evaluation involving a total of 100 participants assessed the systems across six dimensions: (i) Ease of Use, (ii) Relevance, (iii) Transparency, (iv) Responsiveness, (v) Accuracy, and (vi) Likelihood of Recommendation. Based on user feedback and our development experience, we documented twelve key lessons learned, highlighting technical, operational, and ethical challenges affecting the reliability and usability of RAG systems in practice.
Index Terms:
Empirical Software Engineering, AI System Lifecycle, Generative AI, RAG, LLMs, System Design, System Implementation, Human Centered Evaluation
This is the author’s preprint version of a paper accepted to the 51st Euromicro Conference Series on Software Engineering and Advanced Applications (SEAA 2025). The final version will be published by IEEE.
I Introduction
Retrieval-Augmented Generation (RAG) has the capability to enhance Large Language Models (LLMs) by retrieving relevant external knowledge, thereby improving the accuracy of Generative AI (GenAI) applications. While GenAI has seen use in software engineering [1], RAG extends its value to broader domains by combining parametric and non-parametric memory, effectively addressing the limitations of static knowledge bases [2]. Recent advances in retrieval-augmented LLMs enable real-time information retrieval, reducing hallucinations and improving response reliability [3].
The foundational work by Lewis et al. [4] established RAG as a standard for tasks like question answering and knowledge retrieval. Early frameworks such as REALM and RAG demonstrated the benefits of combining dense retrieval with language generation for open-domain tasks [5]. However, most later research has focused on improving retrieval architectures and reducing hallucinations, often evaluated only on clean, English-centric benchmarks [6]. There remains limited exploration of domain-specific, multilingual, or real-world deployments, which this paper addresses through the development and evaluation of RAG systems across diverse application settings, with a focus on retrieval quality and system design.
Accurate information access is important in domains like governance, cybersecurity, agriculture, industrial research, and healthcare. As industries adopt AI for complex tasks, traditional search methods often fall short, especially with multilingual, up-to-date, and contextually relevant knowledge. To address this, we developed RAG systems in collaboration with five organizations: the City of Kankaanpää [^1] https://www.kankaanpaa.fi/, Disarm [^2] https://www.disarm.foundation/framework, AgriHubi [^3] https://maaseutuverkosto.fi/en/agrihubi/, FEMMa [^4] https://www.tuni.fi/en/research/future-electrified-mobile-machines-femma, and a clinical diagnostics group [^5] https://tampere.neurocenterfinland.fi/. Each collaboration guided the system design to meet distinct operational and information access challenges.
This study investigates how RAG systems can be engineered and evaluated in real-world contexts. It emphasizes system design and development for domain-specific applications, user evaluation across criteria such as ease of use and relevance, and lessons that inform future engineering practices. Guided by these objectives, this paper addresses the following three Research Questions (RQs):
  • RQ1: How can RAG systems be designed and developed to address real-world system needs across diverse application domains?
  • RQ2: How do users evaluate domain-specific RAG systems in terms of ease of use, relevance, transparency, responsiveness, and accuracy in real-world applications?
  • RQ3: What are the lessons learned from engineering RAG systems for real-world applications?
To address these research questions, we developed five domain-specific RAG systems and evaluated them through a web-based user study with 100 participants. The evaluation focused on usability, retrieval relevance, transparency, and other user-centered factors. Full methodological details are provided in Section III.
The contributions of this paper are as follows:
  • End-to-end development and deployment of RAG systems for multilingual, domain-specific applications.
  • User-centered evaluation demonstrating real-world performance across usability and accuracy metrics.
  • Practical engineering insights to guide the design of reliable and maintainable RAG pipelines.
  • System-level considerations for integrating RAG into real-world AI-based software, contributing to software engineering practice.
Paper Structure: Section II reviews related work on RAG and its applications. Section III describes the study design. Section IV explains the implementation of five real-world RAG systems. Section V presents user system evaluation, and Section VI outlines key lessons learned. Section VII discusses study limitations, and Section VIII concludes with future directions.
II Related Work
Retrieval-Augmented Generation (RAG) improves the factual accuracy and contextual relevance of Large Language Models (LLMs) by incorporating real-time external information, making it especially valuable for complex tasks such as question answering, legal reasoning, and summarization [7]. Recent work has demonstrated RAG’s utility in taxonomy-driven dataset design [8], token-efficient document handling [7], and multimodal applications that combine text and images via Vision-Language Models (VLMs) such as VISRAG [9]. Despite these advances, OCR noise remains a limiting factor in retrieval fidelity [10]. Ongoing research addresses this by refining dataset construction [8], tackling architectural scalability [11], improving query-document alignment through prompt engineering [12], and applying speculative retrieval to boost performance in multimodal settings [13].
RAG has been widely applied in software engineering to support code understanding and developer tasks. StackRAG [14] leverages Stack Overflow content to enhance developer assistance, while CodeQA [15] employs LLM agents with retrieval augmentation for programming queries. Ask-EDA [16] addresses hallucination reduction in Electronic Design Automation (EDA) via hybrid retrieval. In industrial contexts, Khan et al. [17] examine PDF-focused retrieval challenges, and Xiaohua et al. [18] propose re-ranking and repacking strategies for pipeline optimization. In healthcare, MEDGPT [19] extracts structured insights from PubMed to support diagnostics, while Path-RAG [20] improves pathology image retrieval with knowledge-guided methods. Alam et al. [21] introduce a multi-agent retriever for radiology reports, enhancing clinical transparency, and Guo et al. [22] present LightRAG, a graph-based retriever that boosts knowledge precision across medical domains.
RAG continues to expand into domains like energy and finance. Gamage et al. [23] propose a multi-agent chatbot for decision support in net-zero energy systems, while HybridRAG [24] combines knowledge graphs with vector search to enhance financial document analysis. AU-RAG by Jang and Li [25] dynamically selects retrieval sources using metadata, improving adaptability across sectors. To address retrieval noise, Zeng et al. [26] integrate contrastive learning and PCA for better knowledge filtering. Barnett et al. [27] identify core RAG weaknesses, including ranking errors and incomplete integration, underscoring the ongoing need for more reliable retrieval strategies.
The rise of autonomous AI agents has further improved Retrieval-Augmented Generation (RAG) by enabling self-directed reasoning, adaptive retrieval, and memory persistence. Wang et al. [28] survey LLM-driven agent architectures, while Liu et al. [29] benchmark multi-turn reasoning through AgentBench. AgentTuning by Zeng et al. [30] enhances instruction tuning for retrieval-based decisions. Singh et al. [31] categorize Agentic RAG into single-agent, multi-agent, and graph-based designs, highlighting dynamic tool use. On the retrieval side, Yan et al. [32] introduce CRAG to reduce hallucinations using confidence-based filtering, and Li et al. [33] improve precision through contrastive in-context learning and focus-mode filtering—strengthening RAG’s reliability in complex scenarios.
Conclusive Summary: While RAG continues to advance, challenges in retrieval accuracy, response reliability, and scalability remain [34]. Although hybrid strategies [24], autonomous agents [31], and correction techniques [32, 33, 26] have been explored, there is still limited evaluation in domain-specific settings. This paper addresses that gap by implementing and assessing five RAG systems across key sectors, offering practical insights into their real-world performance and deployment potential.
III Study Design
Figure 1 provides an overview of the methodological steps, from pipeline development to user-centered evaluation.
Figure 1: Overview of the research methodology
To investigate the design and real-world performance of RAG-based systems, we implemented five optimized pipelines across distinct domains and conducted a structured user evaluation to assess their effectiveness and user reception.
III-A Implementing RAG Systems
This section describes how we designed, and built the RAG systems featured in this study. It explains the overall system design, how we selected the case study domains, the unique challenges each domain presented, and the setup used for evaluation.
III-A1 Domain Selection
We selected the application domains to test RAG systems in real-world, knowledge-heavy environments where accurate information retrieval, contextual understanding, and timely decision-making are important. These domains were chosen because they involve different information and require careful decision-making, providing a solid basis to evaluate how well RAG systems can adapt and perform in different settings.
In this study, we apply RAG across five domains: municipal governance, cybersecurity, agriculture, industrial research, and medical diagnostics, to explore how RAG-based retrieval can address diverse domain-specific information needs and support real-world decision-making processes.
III-A2 System Design
Figure 2: Design overview of the RAG systems developed for five domain-specific applications.
The design of the RAG systems in this study follows a two-phase approach:
  • Retrieval Phase: User queries are embedded using pre-trained models (e.g., text-embedding-ada-002) and matched with relevant text chunks via similarity search in vector databases.
  • Generation Phase: The retrieved text chunks are concatenated with the original user query and passed into a large language model (LLM), such as GPT-4o, LLaMA 2 Uncensored, or Poro-34B, to synthesize contextually relevant responses.
This approach improves factual accuracy, minimizes hallucinations, and delivers insights that are well aligned with domain-specific needs.
III-A3 Core Components
Each RAG-based system comprises multiple core components:
  • Data Sources: Knowledge bases include structured and unstructured documents, such as websites, municipal records, cybersecurity reports, agricultural research papers, engineering documents, and clinical guidelines.
  • Vector Database: The retrieved knowledge is stored as vector embeddings in FAISS, Pinecone, or OpenAI’s Vector Store, depending on the system’s latency and scalability requirements.
  • Query Processing: User queries undergo tokenization, embedding generation, and similarity search before being passed to an LLM for the response.
  • Preprocessing Pipelines: Systems rely on PyMuPDF and Tesseract OCR to extract text from PDFs and scanned documents, ensuring the inclusion of both text-based and image-based content. Additionally, for web scraping, the pipeline utilizes BeautifulSoup, Scrapy, and Selenium to extract, clean, and structure data from dynamic and static web pages.
These components enable efficient retrieval and context-aware responses in domain-specific RAG systems.
III-B System Evaluation Method
To understand how the RAG-based systems performed in real usage scenarios, we conducted a structured web-based user study with 100 participants. Each participant was given access to live demo environments and interacted with one or more of the five systems using realistic, domain-specific tasks.
After using the systems, participants completed a standardized survey covering six criteria: Ease of Use, Relevance of Info, Transparency, System Responsiveness, Accuracy of Answers, and Recommendation. The survey included both Likert-scale questions and open-ended feedback. This approach provided both quantitative ratings and qualitative insights into system performance. We reviewed the open-ended feedback to identify common themes in participants’ experiences. We also referred to development notes taken throughout the project. These helped us recognize recurring issues and informed the lessons described in Section VI.
IV Systems Implementation
This section outlines the end-to-end implementation of five RAG-based systems designed for real-world deployment across diverse domains. Each system was developed to address domain-specific retrieval challenges by integrating embedding models, vector databases, and LLMs. The implementations varied based on task complexity, document type, language requirements, and deployment constraints, demonstrating the adaptability of RAG pipelines in applied settings.
  1. 1.
    Kankaanpää City AI: This system enhances transparency of government records. It processes over 1,000 PDFs from 2023–2024, indexing them in FAISS for accurate retrieval of policy documents. The system uses text-embedding-ada-002 as the embedding model to convert documents into vector representations, and gpt-4o-mini as the LLM to generate context-aware responses. This setup allows users to search and access municipal decisions, infrastructure projects, and public policies with ease.
  2. 2.
    Disarm RAG: It is designed to deliver real-time insights into cyber threats, and forensic investigations. It is hosted on a secure server at CSC [^6] https://research.csc.fi/cloud-computing/ (Finnish IT Center for Science), ensuring full data privacy, and uses LLaMA 2-uncensored via Ollama to enable open access to cybersecurity knowledge. The system integrates red team techniques (e.g., phishing, deepfake disinformation, privilege escalation) and blue team strategies (e.g., bot detection, misinformation control, network forensics), grounded in the Disarm Framework. It supports queries such as “How would you create a deepfake to discredit a public figure?” and “What are the latest techniques for bypassing multi-factor authentication (MFA)?”, as well as defensive questions like “How would you detect a disinformation campaign early on?” and “What are effective countermeasures against deepfake-based phishing attacks?”.
  3. 3.
    AgriHubi AI Assist: AgriHubi bridges agricultural policy and practice by processing 200+ Finnish-language PDFs using multilingual OCR and embedding the content into a FAISS vector database. It leverages the Finnish-optimized Poro-34B language model to deliver contextually relevant responses on topics like sustainable farming and soil conservation. The system features a Streamlit chat interface, logs interactions via SQLite, and includes a feedback mechanism for continuous improvement, making agricultural knowledge more accessible to farmers and researchers.
  4. 4.
    FEMMa Oracle: This system optimizes knowledge retrieval for engineering research, particularly in electrified mobile machinery. It processes around 28 PDFs regarding electrified mobile machinery. It integrates GPT-4o and text-embedding-3-large with OpenAI’s Vector Store to enable rapid retrieval of structured engineering research documents. The system ensures that researchers can efficiently access validated technical documentation and structured project information, improving efficiency in engineering-related knowledge retrieval.
  5. 5.
    Assist Doctor: It is an aneurysm diagnostic RAG based application, developed at Tampere University for use by neurologists, radiologists, and vascular surgeons. It retrieves insights from peer-reviewed literature and clinical data using an embedding-based search pipeline and delivers context-aware responses via OpenAI’s GPT-4. With a Streamlit interface, it enables clinicians to access diagnostic criteria, risk stratification models, and treatment comparisons, supporting informed decisions in aneurysm care.
All systems developed in this study comply with GDPR standards to ensure responsible handling of user interactions and system outputs. To support transparency, most systems display source references alongside AI-generated responses. An exception is Disarm RAG, where source citations are omitted due to cybersecurity sensitivity.
V Systems Evaluation
Understanding the real-world effectiveness of RAG-based systems requires moving beyond technical benchmarks to incorporate user-centered evaluation. We conducted a structured user study across five domain-specific deployments, capturing both system performance metrics and user perceptions of trust, relevance, and usability. This practical feedback offers a grounded view of system behavior in real settings and highlights opportunities for targeted improvements.
V-A Participant Demographics and RAG Orientation
To contextualize the system evaluation, we collected detailed background information from the 100 participants involved in the study. Figure 3 illustrates five key dimensions of participant orientation relevant to domain-specific RAG systems: professional role, AI vs. manual search preference, familiarity with RAG, prior usage experience, and comfort with AI-generated outputs.
Figure 3: Participant profiles and their interaction with RAG systems.
Figure 4: User ratings of five RAG systems across six evaluation criteria.
  1. 1.
    Role Distribution: Participants represented five distinct professional categories aligned with our target application domains. Researchers comprised the largest segment (44%), followed by students (20%), domain experts (17%), AI/ML practitioners (16%), and others (3%). This composition reflects a balanced blend of technical stakeholders and domain users, ensuring that the evaluation captures both system-level performance and practical applicability across real-world contexts.
  2. 2.
    AI-Generated vs. Manual Document Search: Participants exhibited a task-sensitive perspective on AI assistance. While a substantial majority (83%) preferred AI-generated responses depending on the nature of the task, only (9%) expressed a consistent preference for AI over manual methods. Conversely, (8%) favored manual search regardless of context. These findings suggest that trust in RAG systems is not absolute but contingent—underscoring the importance of response relevance, transparency, and alignment with user intent.
  3. 3.
    Familiarity with AI-Based RAG: Participants demonstrated a strong familiarity with RAG technologies in general, with (75%) identifying as either moderately (41%) or very familiar (34%) with AI-based RAG systems. However, since many participants were not domain experts in the specific fields covered by the systems (e.g., healthcare, cybersecurity), their feedback primarily reflects their interaction experience with RAG rather than deep subject-matter validation.
  4. 4.
    Experience with AI-Based RAG: Participant engagement with RAG systems was notably high. A majority (82%) reported using such systems either occasionally (45%) or frequently (37%), while only (5%) indicated no prior experience. This distribution reinforces the reliability of the feedback collected, as most evaluations were informed by direct, hands-on interaction rather than hypothetical exposure.
  5. 5.
    Comfort with AI-Generated Responses: Overall, participants expressed high confidence in AI-generated outputs. Nearly three-quarters (73%) reported feeling either mostly (47%) or very comfortable (26%) relying on such responses. Only a small minority (7%) expressed discomfort, indicating a strong baseline of user trust and an encouraging signal for broader adoption of generative AI in domain-specific tasks.
V-B Survey Instrument and Case-wise Findings
Figure 4 presents the aggregated user ratings across six evaluation criteria for all five RAG systems, offering a comparative perspective on system performance.
To capture both measurable and descriptive insights, we employed a survey combining Likert-scale questions (1–5 scale) with open-ended prompts for qualitative feedback. The evaluation focused on the following six core dimensions:
  • Ease of Use: How easy was it to use the system?
  • Relevance of Information: Did the system retrieve relevant and useful information for your queries?
  • Transparency: Did the system show where the information came from?
  • System Responsiveness: How would you rate the system’s responsiveness in retrieving answers?
  • Accuracy of Answers: Based on your knowledge, how accurate were the AI-generated answers provided by the system?
  • Recommendation: Would you recommend this tool to colleagues in your field?
All five RAG systems were evaluated using the same six criteria by a total of 100 participants. The summaries below reflect how each system performed, highlighting key strengths and areas for improvement.
  1. 1.
    Kankaanpää City AI (22 participants): The system performed well in Ease of Use, with (81.8%) rating it as “easy” or “very easy.” Relevance of Info around (82%) and Accuracy of Answers around (91%) were also strong. Transparency was mixed, (45.5%) found it clear, while another (45.5%) found it unclear. (63.6%) said they would recommend the system, suggesting it may be useful in public governance contexts.
  2. 2.
    Disarm RAG (20 participants): Participants reported positive ratings for Ease of Use (65%) and System Responsiveness (75%), despite the complexity of the cybersecurity domain. Relevance of Info and Accuracy of Answers received moderate ratings, while Transparency was low due to intentionally hidden sources. Nevertheless, (55%) of participants indicated they would recommend the system.
  3. 3.
    AgriHubi AI Assist (20 participants): Tailored for Finnish-language agricultural content, the system received strong ratings for Ease of Use (80%) and Accuracy of Answers (65%). Relevance of Info was generally positive, while System Responsiveness and Transparency showed mixed results. Still, (60%) of users responded positively on the Recommendation dimension.
  4. 4.
    FEMMa Oracle (17 participants): The system performed well across all criteria. Accuracy was rated “accurate” or “highly accurate” by (64.7%), and Ease of Use by (82.3%). Relevance of Info was high (88.3%), and (88.9%) found it transparent. Responsiveness was rated “fast” by (50%) and “average” by (28.6%). Overall, (58.8%) said they would recommend it.
  5. 5.
    Assist Doctor (21 participants): Participants found the system easy to use, with (66.7%) rating it as “easy” or “very easy.” Both Accuracy of Answers and Relevance of Info received favourable ratings, each at approximately (62%). System Responsiveness was positively reviewed by more than half of the users. About (62%) found the system transparent, and (47.6%) said they would recommend it.
Across the five systems, Ease of Use and Accuracy of Answers were consistently rated positively. Transparency and Recommendation showed more variation, sometimes due to design choices. For example, Disarm RAG used hidden source information. These differences show that user perception depends on the domain and output presentation.
VI Lessons Learned
While developing and evaluating the five RAG systems, we encountered several technical, operational, and ethical challenges. Based on what we observed during implementation and what participants shared in the evaluation, we summarized a set of lessons that reflect the most common and recurring issues across domains.
VI-A Technical Development
Building RAG systems for real-world applications surfaced a number of technical hurdles that required hands-on problem solving and thoughtful design decisions.
  • Domain-Specific Models Are Essential: General-purpose models like GPT-4o struggled with domain-specific and Finnish-language queries. Leveraging Finnish-optimized models like Poro-34B, along with compatible embedding models (e.g., text-embedding-ada-002), led to more contextually relevant responses.
  • OCR Errors Impact the Pipeline: Noisy OCR output from agriculture and healthcare PDFs degraded FAISS quality. Using TesseractOCR, easyOCR, and regex-based cleanup improved extracted text.
  • Chunking Balances Speed and Accuracy: Token chunk sizes between 200–500 struck a practical balance between retrieval relevance and query latency. Smaller chunks bloated the index, increasing lookup times.
  • FAISS Scalability Hits Limits: With large corpora (>>>10k embeddings), FAISS latency increased noticeably. Metadata filtering by document type reduced search time.
  • Manual Environment Management: Without containerisation, we faced version conflicts across PyTorch, FAISS, OCR libraries, and OpenAI APIs. Strict environment pinning and manual sync across development/production was necessary for stability.
VI-B Operational Factors
Operating RAG systems in real-world settings revealed practical challenges related to data workflows, infrastructure choices, and user interaction management.
  • SQLite for Tracking User Interaction: We used SQLite to log user questions, responses, and ratings (e.g., in AgriHubi). This lightweight store helped identify system failures and understand user behavior.
  • Scraping Pipelines Are Fragile: Websites changed often, breaking parsers. Without stable APIs, we relied on semi-structured feeds and regular script maintenance.
  • Self-Hosted Setup for Speed and Compliance: We hosted LLMs and vector stores on our own servers to reduce GDPR risks and improve speed. This approach balanced control with performance in sensitive domains.
  • Clean Data Boosts Retrieval Quality: Removing OCR noise and duplicates from source data improved answer relevance without modifying models.
  • User Feedback Drives System Tuning: User ratings and comments exposed weak spots, guiding adjustments to retrieval settings and chunk sizes.
VI-C Ethical Considerations
While technical and operational aspects were central to system performance, ethical considerations around transparency, and data bias proved equally important during deployment.
  • Source File References Build Trust: Providing filenames and download links helped users validate AI outputs. In security use cases (e.g., Disarm RAG), sources were intentionally hidden to protect sensitive material.
  • Dataset Bias Impacts Retrieval Balance: Unbalanced source data led to over-representation of some document types. Re-ranking improved diversity and fairness in answers.
Practical and Research Takeaways: Our findings highlight both persistent and emerging challenges in applying RAG systems to real-world, multilingual, and domain-specific settings. While issues like OCR noise, chunk size tuning, and retrieval balancing are well recognized, this study emphasizes the importance of practical strategies such as data cleaning, user feedback mechanisms, and lightweight response validation for improving retrieval quality and system reliability. These lessons extend current research by connecting it to deployment realities and offer value to the software engineering community by addressing concerns related to retrieval infrastructure, stability of data pipelines, and transparency in system outputs. These takeaways help guide the development of adaptable and trustworthy RAG solutions.
VII Study Limitations
This study presents findings grounded in the design, deployment, and evaluation of five domain-specific RAG systems, but several limitations must be acknowledged. First, while our evaluation involved 100 participants across diverse roles including researchers, practitioners, and domain experts, approximately 20% of the sample consisted of students. Although these students had relevant technical or domain experience, their feedback may reflect differing expectations or usage behaviour compared to full-time professionals. This demographic distribution, while broad, could influence the generalizability of findings to strictly industrial settings.
Second, participants interacted with one or more systems, and survey responses were collected separately after each system use. Not all 100 participants engaged with every system; the number of responses per system varied based on individual interest and domain familiarity. For instance, feedback on AgriHubi AI Assist reflects only the users who selected and interacted with that system. This variation in exposure may affect the comparability of results across different systems, and the limited interaction time restricted analysis of longer-term user engagement.
Third, the lessons learned presented in this paper are based on our development experience and observations during system implementation and evaluation. While they do not result from formal empirical analysis, they reflect recurring challenges and design considerations encountered across multiple domains. Although not statistically validated, these insights can inform future work on the design and implementation of RAG systems in applied settings.
VIII Conclusion
In this paper, we presented a tool-assisted approach for designing, implementing, and evaluating RAG-based systems across five real-world domains. Each system was tailored to its specific context—ranging from municipal governance to agriculture and healthcare by integrating multilingual OCR pipelines, semantic retrieval with vector embeddings, and either in-house or cloud-based LLMs. Our user study, involving 100 participants, provided insights into how these systems perform in practice, not just in terms of technical metrics, but also usability, transparency, and user trust.
Through our development work, we identified twelve lessons learned that highlight, in our view, recurring challenges in building practical RAG pipelines. These include balancing chunk size with latency, managing dependencies without containerization, and maintaining retrieval speed at scale. We also found that clean data, user feedback, and clear information presentation are critical for building trust. As industry and research interest in RAG systems grows [5, 34], we hope these insights support future development efforts.
Looking ahead, we see a strong need for more structured evaluation mechanisms that go beyond user ratings. As future work, we propose integrating an Evaluation Agent Model, a system-internal module that checks AI-generated responses for accuracy, relevance, and completeness before presenting them to users. Based on our experiences, user feedback alone is often insufficient to catch factual errors or incomplete responses, especially in domains where missing or misleading information could have serious consequences. An automated evaluation agent could trigger second-stage retrievals or prompt reformulations when weaknesses are detected, creating an adaptive feedback loop. We believe that such mechanisms are essential to improving the reliability and trustworthiness of RAG systems in high-stakes, real-world applications.